OpenStudy (anonymous):

PLEASE HELP Flint purchased a boat for $10,815. He made a down payment of $1,175. He applied for a six-year installment loan with an interest rate of 9.4% in the amount of $9,640. What is the total cost of the boat after six years? $12,649.68 $13,824.68 $11,831.61 $10,546.16

OpenStudy (here_to_help15):

assuming monthly compounding, interest rate = 9.4/1200 = 47/6000 discount rate = 1/(1 +47/6000) = 6000/6047 = D(say) period = 12*6 = 72 let monthly payments be R, then 9640 = RD + RD^2 + RD^3 + ..... RD^72, a GP with a = RD, r = D, sum = RD(1-D^72)/(1-D) plugging in, R = 9640*(47/6047)/((6000/6047)*(1- (6000/6047)^72)) = $175.69 so total amount paid = 72*175.69 + 1175 = $13,824.68 [ B ] <-----

OpenStudy (anonymous):

$13,824.68

OpenStudy (anonymous):

thanks, both of you!